20. What applied horizontal force is required to accelerate a 5 kg along a horizontal surface. With an acceleration of 2 m/s2 if the coefficient of friction is 0.15.


1
Expert's answer
2021-09-29T09:56:07-0400

We can find this force applying Newton's second law:


"ma=F-f, \\
F=ma+f=ma+\mu mg=17.4\text{ N}."